# Squeezed Quasinormal Modes from Nonlinear Gravitational Effects

### v1

@@ II A simple theoretical model to describe quantum effects in quasinormal modes
The source term $S(x,t)$ relevant for second-order (quadratic) quasinormal modes typically includes, up to derivatives, a product of first-order quasinormal modes. The exact regularized form of the source term can be found elsewhere Ioka and Nakano (2007), which also respects the gauge invariant prescription from Garat and Price (2000). Also see Refs. Bucciotti et al. (2024); Redondo-Yuste et al. (2024). Given this, our first objective is to cast this general form of wave equation into a Hamiltonian formalism for the temporal modes, and subsequently estimate the degree of squeezing it generates on the fundamental mode. To estimate this, it is first instructive to simplify the wave equation to see how a quantum mechanical description can be applied to the early-time behavior of long-lived quasinormal modes.
We first consider the solution for the homogeneous part of the above equation using the method of separation of variables. The equation for the radial part looks like a time-independent Schrodinger equation,
$\psi(x,t)\rightarrow\tilde{\psi}(x,\omega):~~\left[-\partial_{x}^{2}+V(x,\omega)\right]\tilde{\psi}(x,\omega)=\omega^{2}\tilde{\psi}(x,\omega),$ (4)
@@
where we have indicated that the potential $V$ could be frequency dependent, $V=V(x,\omega)$, to relate to the quasinormal modes of Kerr black holes Chandrasekhar and Detweiler (1997); Berti et al. (2025); Motohashi (2025). The solutions $\tilde{\psi}(x,\omega)$ typically satisfy the Siegert boundary conditions Siegert (1939): $\tilde{\psi}(x,\omega)\rightarrow e^{\pm i\omega x}$ as $x\rightarrow\pm\infty$ [in the choice of coordinates, $x\rightarrow$-$\infty$ corresponds to the event horizon]. As is well known, such boundary conditions for potentials relevant to the gravitational context imply that the mode frequencies $\omega=\{\omega_{n}\}$ [for given $l,m$, or generically $\omega=\{\omega_{lmn}\}$] are both complex and discrete, with negative imaginary parts. The corresponding solutions, or quasi-normal modes, are well characterized and understood Zerilli (1970); Regge and Wheeler (1957); Maggiore (2008); Berti et al. (2009); Vishveshwara (1970); Teukolsky (1972); Berti et al. (2025), using exact methods or with Wentzel-Kramers-Brillouin (WKB) type approximations Schutz and Will (1985). Once solutions to Eq. (3) are obtained, metric perturbations $h_{\mu\nu}^{(i)},~~i=1,2$ can be reconstructed by following a standardized prescription Berti et al. (2009); Motohashi (2025), while the temporal mode profiles are determined by the discrete quasinormal mode frequencies. In what follows, we suppress the $l,m$ indices often for simplicity, but unless specified otherwise, in the Schwarzschild context, they refer to the $l=4,m=4$ leading second-order quasinormal modes sourced by the $l=2,m=2$ modes.
For the homogeneous part, knowing the solutions of the form,
$h(t)\sim e^{-i\omega_{n}t},~~\text{where}~~\omega_{n}=\pm\omega_{R,n}-i\omega_{I,n},$ (5)
we can also recast the homogeneous part of equation (7) in the following form,
@@
where $\Omega_{n}^{2}=\omega_{R,n}^{2}+\omega_{I,n}^{2}$ and $\omega_{I,n}\geq 0$ in our notation. Here it is evident that the temporal domain of the homogeneous part of the wave equation describes damped quantum harmonic oscillators, with damping coefficients $\gamma_{n}=2\omega_{n,I}$ and proper frequencies $\Omega_{n}=\sqrt{\omega_{R,n}^{2}+\omega_{I,n}^{2}}$. An insightful discussion of this correspondence can be found in Ref. Maggiore (2008).
Since it is the temporal modes that we quantize eventually, it is instructive to recast the original wave equation, Eq. (3), in its mode-expanded form where the spatial modes satisfy the quasinormal mode boundary conditions,
$\sum_{n}\tilde{\psi}_{n}(x,\omega_{n})\left[\ddot{h}_{n}^{(p)}(t)+\omega_{n}^{2}h_{n}^{(p)}(t)\right]=S(x,t),$ (7)
which is an abbreviation for,
$\sum_{n}\tilde{\psi}_{lmn}(x,\omega_{lmn})\left[\ddot{h}_{lmn}^{(p)}(t)+\omega_{lmn}^{2}h_{lmn}^{(p)}(t)\right]=S_{lm}(x,t),$ (8)
after silencing $l,m$ indices. The superscript $(p)$ indicates that we are interested in the particular solutions that account for the presence of a source term. In Eqns. (7)–(8), we have expanded the left-hand side in terms of the known spatial solutions of the homogeneous part that satisfy the quasinormal mode boundary conditions, in the spirit of treating the source term as effective force perturbations on the known homogeneous solutions. The temporal parts describe quantum harmonic oscillators, but with complex frequencies. We are interested in scenarios where the real part of the frequency dominates substantially, such that the quasinormal modes are long-lived. Practically, these are the modes that dominate the ringdown signal. The quantum behavior of these long-lived modes will then be effectively captured by the dynamics of their real frequency part.
II.1 Projection to quasinormal modes
@@
To pick out the response of individual quasinormal modes, we need to consider some type of orthogonality relations for quasinormal modes. This is not straightforward, because the spatial modes of quasinormal modes are not square-integrable. Several alternatives that respect the “quasinormal” aspect of the modes have been considered in the literature Ching et al. (1995); Leung et al. (1997, 1998); Jaramillo et al. (2021); Green et al. (2023); London and Gurevich (2024); Motohashi (2025); Sauvan et al. (2022); Leaver (1986); Kristensen et al. (2020). They explore different approaches including include the use of Green’s functions Leaver (1986); Ching et al. (1995) and bi-orthogonal bases. Here one finds states $\{|\phi_{j}\rangle\}$ that are orthogonal to $\{|\tilde{\psi}_{n}\rangle\}$ such that the inner product $\langle\phi_{j}|\tilde{\psi}_{n}\rangle\propto\delta_{jn}$. Electromagnetic/optical analogies are discussed in Refs. Kristensen et al. (2020); Leung et al. (1994).
@@
Our choice of inner product is motivated by treatment in Refs. Leung et al. (1994); Ching et al. (1995); Motohashi (2025); Leung et al. (1997). To define the inner product, we begin with the abbreviated notation $\psi_{j}=\tilde{\psi}(\omega_{j},x).$ It follows from Eq. (4) that,
@@
when $n\neq j$. The limit $\omega_{k}\rightarrow\omega_{j}$ can be taken continuously along an appropriate contour in the complex plane of quasinormal mode frequencies, exploiting the circumstance that the frequencies $\{\omega_{n}\}$, although discrete for a given black hole, depend on continuously variable black hole parameters such as the black hole mass and spin. With the identification, $H=[-\partial_{x}^{2}+V(x,\omega)]$, we see from Eq. (4) that $H\psi_{j}=\omega_{j}^{2}\psi_{j}.$ This allows us to define the following operator (where the operations are performed in the order they appear),
@@
where the norm $\mathcal{N}_{j}^{2}$ is complex in general Motohashi (2025); Leung et al. (1994); Ching et al. (1995). The action of $\mathcal{L}(\psi_{j})[\cdot]$ on the left-hand side of Eq. (7) only retains the coefficient of the $j^{\text{th}}$ mode in the expansion, with the normalization factor $\mathcal{N}_{j}^{2}$. The action of $\mathcal{L}(\psi_{j})[\cdot]$ on the right-hand side of Eq. (6) must therefore determine the driving term of the wave equation with appropriate pre-factors. Hence, we can write,
$\left[\ddot{h}_{j}^{(p)}(t)+\omega_{j}^{2}h_{j}^{(p)}(t)\right]=\mathcal{N}_{j}^{-2}\mathcal{L}(\psi_{j})[S(x,t)]=g_{j}(t).$ (20)
@@
In the gravitational context, we know that the temporal part of the source contains the product of lower-order quasinormal modes, dominated by the fundamental ($\omega_{220}\equiv\omega_{0}$) mode Ioka and Nakano (2007). More generically, we can approximate, to the leading order, $g_{j}(t)\approx\sqrt{2}\Lambda_{j}[h_{j,0}(t)]^{2}$, where we have denoted $h_{j,0}$ as the fundamental mode having exactly half the frequency to that of the second harmonic mode $h_{j}^{(p)}(t)$. Recall that here the superscript $(p)$ denotes that $h_{j}^{(p)}(t)$ is the particular solution to Eq. (20). We have also absorbed any pre-factors to the effective parameter $\Lambda_{j}$, determined by Eq. (20). While these results apply to any of the modes, we restrict ourselves to the long-lived modes that have a substantial real frequency (propagating) part, which are also the ones likely to be observed. We thus obtain,
$\ddot{h}_{j}^{(p)}(t)+\omega_{j}^{2}h_{j}^{(p)}(t)\approx\sqrt{2}\Lambda_{j}[h_{j,0}(t)]^{2},$ (21)
@@
where $h_{j,0}$ satisfies $\ddot{h}_{j,0}+\omega_{j,0}^{2}h_{j,0}=0$. The equation resembles the equation for a forced harmonic oscillator, except note that $\{\omega_{j}\}$ and potentially $\Lambda_{j}$, are complex. By considering long-lived modes, to the leading order, we may treat these parameters to be real and neglect the contributions from their imaginary parts. This only applies to lower-order quasinormal modes, as the higher-order ones are increasingly damped; however, the lower-order modes are also the ones of interest here.
@@
Re-instating the notations in the Schwarzchild context as an example, to the leading order ($j=0$, $\Lambda_{0}\equiv\Lambda$), we have,
$\ddot{h}_{44}^{(p)}(t)+\omega_{440}^{2}h_{44}^{(p)}(t)\approx\sqrt{2}\Lambda[h_{220}(t)]^{2}~~\text{and}~~\ddot{h}_{kk0}(t)+\omega_{kk0}^{2}h_{kk0}(t)=0,~~k\in\{2,4\}.$ (22)
With $k=4$ in the second equation is of course the solution to the homogeneous part of the first equation. Also note that Eq. (22) is a slight generalization of the standard equation we often see in higher harmonic generation, since the source term has a frequency, $\omega_{220}\neq\omega_{440}$. Assuming a solution $h_{220}\approx a\cos{(\omega_{220}t)}$, for the fundamental mode, we obtain that the particular solution to the above equation has the following form,
$h_{44}^{(p)}(t)=\frac{a^{2}\Lambda}{\sqrt{2}\omega_{440}^{2}}+\frac{a^{2}\Lambda\cos(2\omega_{220}t)}{\sqrt{2}(\omega_{440}^{2}-4\omega_{220}^{2})}.$ (23)
Note that the temporal part of the particular solution oscillates at frequency $\omega_{44}^{(p)}=2\omega_{220}$, which is the second harmonic generated mode we are interested in. Anticipating quantization, we extract the Hamiltonian corresponding to the driving force,
$H_{I}\approx-\sqrt{2}\Lambda\left[\hat{h}_{220}(t)\right]^{2}\hat{h}_{44}^{(p)}(t).$ (24)
Returning to more general notation, let us consider the creation and annihilation operators for the field in the temporal domain $A_{\omega}$ for the fundamental mode and $A_{\omega}^{(p)}\equiv A_{\omega_{p}}\equiv A_{2\omega}$ for the second harmonic mode, we can rewrite the interaction Hamiltonian as
$H_{I}\approx-\frac{\hbar\lambda}{2}\left[(A_{\omega}^{\dagger})^{2}e^{-2i\omega_{0}t}+A_{\omega}^{2}e^{2i\omega_{0}t}+2A_{\omega}^{\dagger}A_{\omega}+1\right]\left(A_{\omega_{p}}e^{i\omega_{p}t}+A_{\omega_{p}}^{\dagger}e^{-i\omega_{p}t}\right),$ (25)
where $\lambda\propto\Lambda\sqrt{\hbar}$ [see Appendix. A where $\lambda$ is estimated for gravitational radiation quantized within a cubical box of reduced volume $V=(c/\omega)^{3}$, where $\omega$ is the frequency of the radiation and $c$ is the speed of light.]. We can make a canonical transformation,
and
so that the interaction Hamiltonian can be represented as,
Note that the fundamental quasinormal mode of the Schwarzschild spacetime is known to have a substantial (although not too large) quality factor $Q\approx 4.2$. While the quality factor remains fixed, the lifetime can increase with the mass of the black hole Berti et al. (2009). They could serve as sources of long-lived quasinormal modes, adequately described by the real-frequency (oscillatory) part over short time-scales relevant to resonant detection. The situation is especially favorable for rotating (Kerr) black holes, which are the ones also observed in LIGO. Here, the quality factors and lifetimes of the fundamental quasinormal modes can be substantially higher. For instance, the quality factors can increase with the black hole spin, approaching very large values near extremal conditions  Berti et al. (2009); Detweiler (1980); Hod (2008), even approaching the zero-damped limit Yang et al. (2013a, b). They correspond to extreme scenarios where our analyses become exact.
II.2 Approach to estimate $\lambda$ based on phenomenology
Above we have reduced the problem to a single effective parameter $\lambda$, potentially evaluated from the source term $S(x,t)$ using Eq. (20), or another appropriate choice of projection of the source term onto the quasinormal mode basis. It may be helpful to consdier a simpler phenomenological model that also allows for estimating the coupling $\lambda$ independently, and in principle, can be applied to extract the coupling from observed amplitudes in rather generic scenarios. To this end, we note that the amplitude $\frac{\lambda}{2}$ is determined by the fraction of energy dissipated via second harmonic generation, which numerical general relativity provides good estimates, that the amplitude of the second harmonic mode can be $\sim 10\%$ to that of the fundamental mode in some scenarios Cheung et al. (2023); Mitman et al. (2023). In fact, we can take the semi-classical limit to treat the fundamental mode in a coherent state $A_{\omega}|\alpha_{\omega}\rangle=\alpha_{\omega}|\alpha_{\omega}\rangle$, and approximate the interaction Hamiltonian as,
It seems then appropriate that we could determine $\lambda$ at short times from simulations or actual data as,
III Degree of Squeezing
The interaction Hamiltonian in Eq. (29) describes second harmonic generation, which is a well-known method in quantum optics to generate and probe sub-Poissonian states of the radiation field Mandel (1982). We can extend the analysis to the gravitational context as well, by considering quasinormal modes that are long-lived, and by restricting to timescales that are much smaller than the lifetime of the quasinormal modes. With this approximation, we can proceed to estimate the degree of squeezing generated in the ringdown phase. To this end, note that the Hamiltonian from Eq. (29) leads to the following set of equations that describe second harmonic generation,

### v2

@@ Appendix A Classical and quantum amplitudes of gravitational radiation
The source term $S(x,t)$ relevant for second-order (quadratic) quasinormal modes typically includes, up to derivatives, a product of first-order quasinormal modes. The exact regularized form of the source term can be found elsewhere Ioka and Nakano (2007), which also respects the gauge invariant prescription from Garat and Price (2000). Also see Refs. Bucciotti et al. (2024); Redondo-Yuste et al. (2024). Given all this, our first objective is to cast this general form of wave equation into a Hamiltonian formalism for the temporal modes. This includes non-linear terms that represent second harmonic generation, which in turn implies squeezing.
Based on the driven quasinormal mode response predicted by Eq. (39), we consider a quasinormal mode expansion for part of the second-order solution in the near-field region to make a leading-order estimate for the coupling. This estimate will be approximate. Quasinormal modes do not form a complete basis, so the expansion is delicate and non-uniform in space-time. At a given time shortly past merger it is expected to be most accurate near the source, where most of the amplitude resides, so that the expansion coefficients can be read off (approximately) from behavior in that region. Careful treatment would require bounding the errors due to spatial cutoff and incompleteness of the basis, and is left for future work. Our treatment, motivated by the standard approaches to obtaining couplings in the context of normal (zero-damped) modes subject to external forcing, assumes a reasonable cutoff (i.e., neglect of the formally infinite far-field contribution), takes initial mode occupations as input and evolves them for a few cycles to establish equilibrium. The resulting estimate is far from rigorous, but it supports the semi-phenomenological estimate and indicates how it could be made more precise.
We begin by considering the solution for the homogeneous part of Eq. (39) using the method of separation of variables. The following candidate solution for the homogeneous part of Eq. (39) can be considered, $\psi(x,t)\sim\int d^{2}\omega\tilde{\psi}(x,\omega)h(t,\omega)$, where $\omega$ is the complex parameter independent of $x,t$ such that $\ddot{h}+\omega^{2}h=0$ for the temporal part, and,
$\left[-\partial_{x}^{2}+V(x,\omega)\right]\tilde{\psi}(x,\omega)=\omega^{2}\tilde{\psi}(x,\omega),$ (40)
@@
for the spatial part, which resembles a time-independent Schrodinger equation. Above, we have indicated that the potential $V$ could be frequency dependent, $V=V(x,\omega)$, to relate to the quasinormal modes of Kerr black holes Chandrasekhar and Detweiler (1997); Berti et al. (2025); Motohashi (2025). Applying the boundary conditions typically pick specific values for the frequencies, $\omega$. The solutions $\tilde{\psi}(x,\omega)$ in the quasinormal mode context have to satisfy the Siegert boundary conditions Siegert (1939): $\tilde{\psi}(x,\omega)\rightarrow e^{\pm i\omega x}$ as $x\rightarrow\pm\infty$ [in the choice of coordinates, $x\rightarrow$-$\infty$ corresponds to the event horizon]. As is well known, such boundary conditions for potentials relevant to the gravitational context imply that the mode frequencies $\omega=\{\omega_{n}\}$ [for given $l,m$, or generically $\omega=\{\omega_{lmn}\}$] are both discrete and complex, with negative imaginary parts. The corresponding solutions, or quasi-normal modes, are well characterized and understood Zerilli (1970); Regge and Wheeler (1957); Maggiore (2008); Berti et al. (2009); Vishveshwara (1970); Teukolsky (1972); Berti et al. (2025), using exact methods or with Wentzel-Kramers-Brillouin (WKB) type approximations Schutz and Will (1985). Once the solutions are obtained, metric perturbations $h_{\mu\nu}^{(i)},~~i=1,2$ can be reconstructed by following a standardized prescription Berti et al. (2009); Motohashi (2025), while the temporal mode profiles are determined by the discrete quasinormal mode frequencies.
To estimate the coupling to leading order perturbatively, we consider a particular ansatz of the form $\psi^{(p)}(x,t)\approx\sum_{n}\tilde{\psi}_{n}(x,\omega_{n})h_{n}^{(p)}(t)$ to describe a part of the second-order solution, which can be understood as the driven response of the first order quasinormal modes to the leading order. Such quasinormal mode expansions are generically not complete, and the applicability is best limited to finite regions of spacetime Ching et al. (1996); Szpak (2004); Settimi et al. (2009). By restricting to near-field and quasinormal mode expansions, we mainly miss out on the asymptotic (far-field) behavior of the source and the response.
The approximations we make here are quite analogous to defining a Fabry-Perot-like cavity of finite length in an optical setting, with open boundary conditions on both ends. Here, with an appropriate definition of an inner product, quasinormal mode expansions are complete within the cavity, and can be used to make meaningful predictions comparable to a normal mode setting Settimi et al. (2009). For readers interested in a complete description of the solution at second order, we recommend that earlier works using Green’s functions Perrone et al. (2024); Andersson (1997); Okuzumi et al. (2008); Szpak (2004); Lagos and Hui (2023); Leaver (1986); Nakano and Ioka (2007); Ioka and Nakano (2007); Ching et al. (1996) are more appropriate.
It is instructive to first consider the limit when the source term reduces to zero to understand the unperturbed sector. Substituting the expansion $\psi(x,t)=\sum_{n}\tilde{\psi}(x,\omega_{n})h_{n}(t)$, the homogeneous part of the wave equation can be expressed as,
$\sum_{n}\tilde{\psi}_{n}(x,\omega_{n})\left[\ddot{h}_{n}(t)+\omega_{n}^{2}h_{n}(t)\right]=0,\$(41)
where we have used Eq. (40). Knowing the solutions for the temporal part of quasinormal modes of the form,
$h_{n}(t)\sim e^{-i\omega_{n}t},~~\text{where}~~\omega_{n}=\pm\omega_{R,n}-i\omega_{I,n},$ (42)
we can also write,
@@
where $\Omega_{n}^{2}=\omega_{R,n}^{2}+\omega_{I,n}^{2}$ and $\omega_{I,n}\geq 0$ in our notation. Here, focusing on individual terms of the sum, it is evident that the temporal domain of the homogeneous part of the wave equation describes damped quantum harmonic oscillators, with damping coefficients $\gamma_{n}=2\omega_{n,I}$ and proper frequencies $\Omega_{n}=\sqrt{\omega_{R,n}^{2}+\omega_{I,n}^{2}}$. An insightful discussion of this correspondence can be found in Ref. Maggiore (2008). Here, it is also evident that the individual quasinormal modes in an effective quantum description could be thought of as damped quantum harmonic oscillators. The effect of incorporating the source term can then be understood as forcing. In the temporal domain, the system can be described as a collection of forced, damped, quantum harmonic oscillators, where, given the temporal dependence of the source term as quadratic in first-order modes, each of the modes in the above sum contributes to the second harmonic mode. To second order, each of these forced harmonic oscillators will be vibrating at the frequency of the forcing term.
B.1 The driven quasinormal mode response
To account for the effect of the source term as small, driving effects on the first-order modes, we decompose the second-order response and the spatial part of the source term $S$, in terms of the first-order mode functions, $\{\tilde{\psi}_{n}\}$, in the near-field region where we aim to estimate the leading order coupling. While this can be achieved in different problems where $\{\psi_{n}\}$ form a complete basis within the region of interest, such a decomposition can only account for parts of the solution and source term “parallel” (in the sense of an appropriate projection) to the quasinormal mode functions since quasinormal modes do not form a complete basis. However, the parts expandable in terms of the mode functions $\{\psi_{n}\}$ in finite regions of interest is expected to capture the driven quasinormal mode response as discussed below.
The part of the source, $S_{\parallel}$, which can be expanded within the region of interest in terms of $\{\psi_{n}\}$ can be identified from the source $S$ by an appropriate projection method, using a projection operator $\mathcal{L}(\tilde{\psi}_{n})[\cdot]$ satisfying $\mathcal{L}(\tilde{\psi}_{n})[\tilde{\psi}_{m}]\propto\delta_{nm}$ for a given region, $\mathcal{C}$. We discuss an exemplary choice for the projection operator in Sec. B.2, valid for any contour $\mathcal{C}$. A norm $\mathcal{N}_{n}^{2}$ can also be defined using the projection operator $\mathcal{L}(\tilde{\psi}_{n})[\cdot]$ as $\mathcal{N}_{n}^{2}:=\mathcal{L}(\tilde{\psi}_{n})[\tilde{\psi}_{n}]$ (also see Ref. Motohashi (2025)). In what follows, we take $\mathcal{C}$ to be the near-field—vicinity of the peak of the potential barrier. Other alternatives for $\mathcal{L}(\tilde{\psi}_{n})[\cdot]$ are also certainly possible, and our prescription here is largely agnostic to the particular choice we make in Sec. B.2. This is because, using such a projection operator within the region of interest, we can write,
$S=S_{\parallel}+S_{\perp},~~\text{where}~~S_{\parallel}:=\sum_{n}\mathcal{N}_{n}^{-2}\mathcal{L}(\tilde{\psi}_{n})[S]\tilde{\psi}_{n}.$ (44)
From Eq. (44), it also follows that, $\mathcal{L}(\tilde{\psi}_{m})[S_{\parallel}]=\mathcal{L}(\tilde{\psi}_{m})[S]$. Above, $S_{\perp}=S-S_{\parallel}$, and it can be noted that $\mathcal{L}(\tilde{\psi}_{n})[S_{\perp}]=0~~\forall n$. It is also evident that $S_{\parallel}$ is the part whose spatial mode profile can be expanded in terms of the first-order modes, $\tilde{\psi}_{n}$, within the region $\mathcal{C}$. Naturally, if $\{\psi_{n}\}$ were to form a complete basis for the problem or the region of interest, $S_{\parallel}=S$ and $S_{\perp}=0$.
In the quasinormal mode context, a complete description, including the tails, must account for the entire source term, and this is studied in Refs. Cheung et al. (2023); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007); Okuzumi et al. (2008); Lagos and Hui (2023); Perrone et al. (2024). Owing to the lack of completeness for the first order quasinormal modes to describe either the second-order source term or the solution, the complete particular solution must be written as $\phi^{(p)}(x,t)=\psi^{(p)}(x,t)+\chi^{(p)}(x,t)$, where $\chi^{(p)}(x,t)$ satisfies,
$\left[\partial_{t}^{2}-\partial_{x}^{2}+V\right]\chi^{(p)}=S_{\perp}.$ (45)
As the driven quasinormal mode response is sourced by $S_{\parallel}$ to the leading order, we can reduce the dynamics to that of driven oscillators from the following wave equation,
$\left[\partial_{t}^{2}-\partial_{x}^{2}+V\right]\psi^{(p)}\approx S_{\parallel}.$ (46)
Here, $\psi^{(p)}(x,t)\approx\sum_{n}\tilde{\psi}_{n}(x,\omega_{n})h_{n}^{(p)}(t)$ and $S_{\parallel}$ also permits a mode-expansion in terms of the first-order quasinormal modes, $\tilde{\psi}_{n}(x,\omega_{n})$. Using the above ansatz, we can rewrite Eq. (46) as,
$\sum_{n}\tilde{\psi}_{n}(x,\omega_{n})\left[\ddot{h}_{n}^{(p)}(t)+\omega_{n}^{2}h_{n}^{(p)}(t)\right]=S_{\parallel}.$ (47)
We have again made use of Eq. (40) in Eq. (47), and the latter is an abbreviation for,
$\sum_{n}\tilde{\psi}_{lmn}(x,\omega_{lmn})\left[\ddot{h}_{lmn}^{(p)}(t)+\omega_{lmn}^{2}h_{lmn}^{(p)}(t)\right]=S_{\parallel,lm},$ (48)
after silencing $l,m$ indices. For our considerations, we restrict to the $l=4,m=4$ leading second-order quasinormal modes sourced by the $l=2,m=2$ mode. This allows us to project out the spatial part on both ends and reduce the equations to that of a forced damped harmonic oscillator in the temporal domain, from which an interaction Hamiltonian can be derived.
The source term in the near-field region to leading order is approximated as $S\approx S_{\parallel}$ – that is, a superposition of the spatial part of the first order modes, $\tilde{\psi}_{n}(x,\omega)$. Physically, this provides an approximate description of the spatial part of the second harmonic in the near-field, as it is produced by forcing the first-order quasinormal modes to the second harmonic.
B.2 Projection to quasinormal modes
@@
To pick out the response of individual quasinormal modes, we need to consider some type of orthogonality relations for quasinormal modes. This is not straightforward, because the spatial modes of quasinormal modes are not square-integrable. Several alternatives that respect the “quasinormal” aspect of the modes have been considered in the literature Ching et al. (1995); Leung et al. (1997, 1998); Jaramillo et al. (2021); Green et al. (2023); London and Gurevich (2024); Motohashi (2025); Sauvan et al. (2022); Leaver (1986); Kristensen et al. (2020). They explore different approaches including include the use of Green’s functions Leaver (1986); Ching et al. (1995) and bi-orthogonal bases. Here one finds states $\{|\phi_{j}\rangle\}$ that are orthogonal to $\{|\tilde{\psi}_{n}\rangle\}$ such that the inner product $\langle\phi_{j}|\tilde{\psi}_{n}\rangle\propto\delta_{jn}$. Electromagnetic/optical analogies are discussed in Refs. Kristensen et al. (2020); Leung et al. (1994); Settimi et al. (2009).
@@
Here we discuss an exemplary choice for the projection operator, which is motivated by treatment in Refs. Leung et al. (1994); Ching et al. (1995); Motohashi (2025); Leung et al. (1997). This by no means is expected to be a unique choice. Alternatives are certainly possible, and our analysis only assumes the availability of such a projection operator for quasinormal modes. To define the projection operator, we begin with the abbreviated notation $\psi_{j}=\tilde{\psi}(\omega_{j},x).$ It follows from Eq. (40) that,
@@
when $n\neq j$. As a formal device, the limit $\omega_{k}\rightarrow\omega_{j}$ can be taken continuously along an appropriate contour in the complex plane of quasinormal mode frequencies, exploiting the circumstance that the frequencies $\{\omega_{n}\}$ (as well as $U(x,\omega_{k})$ and $\psi_{k}$), although discrete for a given black hole, depend on continuously variable black hole parameters such as the black hole mass and spin. Hence, the intermediate values of $\omega_{k}$ are also physically meaningful (also see Ref. Motohashi (2025)). With the identification, $H=[-\partial_{x}^{2}+V(x,\omega)]$, we see from Eq. (40) that $H\psi_{j}=\omega_{j}^{2}\psi_{j}.$ This allows us to define the following operator (where the operations are performed in the order they appear),
@@
where the norm $\mathcal{N}_{j}^{2}$ is complex in general Motohashi (2025); Leung et al. (1994); Ching et al. (1995).
B.3 Estimating the driving force
@@
The action of an appropriate projection operator, $\mathcal{L}(\psi_{j})[\cdot]$, on the left-hand side of Eq. (47) only retains the coefficient of the $j^{\text{th}}$ mode in the expansion, with the normalization factor $\mathcal{N}_{j}^{2}$. The action of $\mathcal{L}(\psi_{j})[\cdot]$ on the right-hand side of Eq. (47) must therefore determine the driving term of the wave equation with appropriate pre-factors. Hence, we can write,
$\left[\ddot{h}_{j}^{(p)}(t)+\omega_{j}^{2}h_{j}^{(p)}(t)\right]=\mathcal{N}_{j}^{-2}\mathcal{L}(\psi_{j})[S_{\parallel}]=\mathcal{N}_{j}^{-2}\mathcal{L}(\psi_{j})[S]=g_{j}(t),$ (60)
@@
where we have also used Eq. (44). In the gravitational context, we know that the temporal part of the source contains the product of lower-order quasinormal modes, dominated by the fundamental ($\omega_{220}\equiv\omega$) mode Ioka and Nakano (2007). More generically, we can approximate, to the leading order, $g_{j}(t)\approx\sqrt{2}\Lambda_{j}[h_{\omega}(t)]^{2}$, where we have denoted $h_{\omega}$ as the fundamental mode having exactly half the frequency to that of the second harmonic mode $h_{j}^{(p)}(t)$. Recall that here the superscript $(p)$ denotes that $h_{j}^{(p)}(t)$ is the particular solution to Eq. (60). We have also absorbed any pre-factors to the effective parameter $\Lambda_{j}$, determined by Eq. (60). While these results apply to any of the modes, we restrict ourselves to the long-lived modes that have a substantial real frequency (propagating) part, which are also the ones likely to be observed. We thus obtain,
$\ddot{h}_{j}^{(p)}(t)+\omega_{j}^{2}h_{j}^{(p)}(t)\approx\sqrt{2}\Lambda_{j}[h_{\omega}(t)]^{2},$ (61)
@@
where $h_{\omega}$ satisfies $\ddot{h}_{\omega}+\omega^{2}h_{\omega}=0$. Multiplying by an effective moment of inertia $\mu_{j}\sim c^{5}/(G\omega_{j}^{3})$ across, the equation resembles the equation for a forced harmonic oscillator, except note that $\{\omega_{j}\}$ and potentially $\Lambda_{j}$, are complex. By considering long-lived modes, to the leading order, we may treat these parameters to be real and neglect the contributions from their imaginary parts. This only applies to lower-order (small $j$) quasinormal modes, as the higher-order ones are increasingly damped; however, the lower-order modes are also the ones of interest here.
@@
Re-instating the notations in the Schwarzchild context as an example, to the leading order, we have,
$\ddot{h}_{44j}^{(p)}(t)+\omega_{44j}^{2}h_{44j}^{(p)}(t)\approx\sqrt{2}\Lambda_{j}[h_{220}(t)]^{2}~~\text{and}~~\ddot{h}_{kkj}(t)+\omega_{kkj}^{2}h_{kkj}(t)=0,~~k\in\{2,4\}.$ (62)
With $k=4$ in the second equation, it is, of course, the homogeneous part of the first equation. This motivates the suggestive form for the coupled wave equation Eq. (4) used in the main text.
Appendix C Exact analysis of squeezing in second harmonic generation
Here we briefly revisit a perturbatively exact treatment to estimate the degree of squeezing in second harmonic generation by closely following the analysis in Refs. Ou (1994); Li and Kumar (1994); Suhara et al. (1996), which the readers may refer to for details. In comparison, our analysis is in the temporal domain. Consider the following dynamical equations that describe second harmonic generation,